### v1

# Scaling Test-time Compute for Low-resource Languages: Multilingual Reasoning in LLMs

Abstract
Recent advances in test-time compute scaling have enabled Large Language Models (LLMs) to tackle deep reasoning tasks by generating a chain-of-thought (CoT) that includes trial and error, backtracking, and intermediate reasoning steps before producing the final answer. However, these techniques have been applied predominantly to popular languages, such as English, leaving reasoning in low-resource languages underexplored and misaligned. In this work, we investigate the multilingual mechanism by which LLMs internally operate in a latent space biased toward their inherently dominant language. To leverage this phenomenon for low-resource languages, we train models to generate the CoT in English while outputting the final response in the target language, given input in the low-resource language. Our experiments demonstrate that this approach, named English-Pivoted CoT Training, outperforms other baselines, including training to generate both the CoT and the final response solely in the target language, with up to 28.33% improvement. Further analysis provides novel insights into the relationships between reasoning and multilinguality of LLMs, prompting for better approaches in developing multilingual large reasoning models.
[…]
123
1 Introduction
Large Language Models (LLMs) have showcased exceptional performance in a wide-range of task [35, 34], particularly in English, due to the vast amount of available data during pre-training [22, 6, 11, 27]. Recent advances in post-training techniques, such as test-time scaling [19], which explicitly trains models to generate chain-of-thought (CoT) reasoning, have significantly enhanced model accuracy in complex reasoning tasks, notably in mathematics and programming problems. However, such advances have unintentionally widened the performance gap between popular languages like English and low-resource languages [10, 5]. This disparity arises primarily because post-training methods typically require high-quality, manually curated datasets, predominantly available only in English. Additionally, cross-lingual misalignment and inherent biases within multilingual training corpora further contribute to the problem.
Figure 1: Illustrative example of model behavior (r1-distill-Llama-8B) when prompted with the same problem in English (robust reasoning) versus a low-resource language - Irish (reduced understanding and conversational ability). Training with an Irish chain-of-thought diverges from baseline, while training with English chain-of-thoughts achieves the best of both worlds.
Figure 1 illustrates this phenomenon by comparing outputs from a small-size state-of-the-art reasoning model (r1-distill-Llama-8B [4]) given the same mathematical problem presented in English and in a low-resource language (Irish). When the question is presented in English, the model reasoning process can lead to the correct solution. In contrast, when the same problem is prompted in Irish, the model misunderstands the problem, unable to converse in the target language, causing the reasoning process to fail and resulting in an incorrect solution. Our proposed solution (shown on the right side of the figure) addresses this issue by aligning the model’s reasoning process across languages, allowing it to “think” internally in its dominant language (English), thereby substantially improving performance while enhancing user experiences in interacting in low-resource languages.
Multilingual reasoning, which combines logical inference with multilingual capabilities, is essential for creating AI systems that can operate effectively across diverse linguistic and cultural contexts. Despite its importance, this remains a relatively unexplored domain, with most existing efforts concentrating on a limited set of high-resource languages [5], leaving low-resource and extremely low-resource languages underrepresented. Current multilingual alignment methods can be split into three broad categories: (1) fine-tuning (including language-alignment [36] or multilingual reasoning fine-tuning [13, 1]), (2) prompting strategies (translation-based [23] or self alignment-based prompting [21]), and (3) model editing techniques (model merging [31, 24], layer swapping [2]). Fine-tuning strategies require large amounts of data in low-resource languages to let the model understand and reason in those languages. Prompt-based methods and model editing approaches rely heavily on the model’s existing comprehension of the low-resource language, which is often insufficient. In general, while these methods show promise, they often require extra training data, additional modules, or incur translation errors, limiting their practical usability in truly low-resource settings.
Inspired by recent advances in test-time scaling, where models explicitly generate reasoning traces without the need for carefully crafted reasoning prompts, and recent findings on the dynamics of language models across layers [23, 30, 26] (where distinct layers/neurons specialize in language understanding and others in reasoning, typically biased towards English), we explore the connections between reasoning and multilinguality, particularly how reasoning can be transferred across languages by keeping the explicit chain-of-thought in English, while boosting the low-resource language understanding and generation of the model while keeping inputs and final responses in the target language. Our method allows the LLM to interpret a problem in the target language, perform the reasoning steps in English, where it is most robust, and then subsequently generate accurate final responses in the target language.
Our contributions:
  • We propose a novel and effective method for adapting an LLM’s reasoning capabilities to low-resource languages at inference time. Our approach explicitly aligns the reasoning process across languages, addressing linguistic misalignment and boosting performances in reasoning problems in low-resource languages.
  • Our results demonstrate successful transfer of test-time scaling to low-resource languages by leveraging the model’s internal reasoning representations across languages and models, upto 28.33% performance improvement.
  • Our analysis provides new insights into the multilinguality of LLMs by explicitly separating language understanding and reasoning processes.
  • We introduce LC2024, the first-ever benchmark dataset for evaluating mathematical reasoning in Irish. This LaTeX-formatted dataset is derived from the Irish Leaving Certificate 2024 mathematics exam.
Source code, model weights, and datasets will be made publicly available for future research and benchmarking purposes.
2 Related Works
2.1 Multilingual Reasoning in LLMs
Reasoning is formally described as the cognitive process of logically analyzing available information to reach conclusions, enabling both humans and AI systems to address complex problems and decisions. [5]. Chain-of-thought [29, 32] has emerged as a powerful technique for improving reasoning in LLMs. By prompting the model to generate explicit step-by-step reasoning, for example, simply instructing the model with “Let’s think step by step”, CoT can significantly increase accuracy on arithmetic, logical, and commonsense problems. CoT prompting was initially studied mostly in English, but recent work shows it can also elicit reasoning in other languages [23, 21], for example, by prompting the model to translate to English first and then solve the task in English. [23] finds that few-shot examples, prompting with chain-of-thought in English, consistently achieve competitive or better performance than reasoning in the native language of the question. To bridge the gap between languages, alignment techniques have been carried out to align representation between low-resource languages and English through learning with parallel sentences [36, 14] or additional multilingual encoders [9]. On the other hand, recent works also attempt to directly fine-tune LLMs in reasoning tasks across multiple languages, by obtaining reasoning data in low-resource languages, mostly through neural machine translation [1, 13].
While prior works highlight stronger performance when aligning the language model to reason in English for tasks in low-resource languages, they typically leverage prompting techniques or require additional modules or data resources. In this work, we explicitly train LLMs to separate their reasoning and language understanding capabilities by separating the languages used in internal reasoning and final response.
2.2 Alignment of Internal Representations in Multilingual LLMs
There has been a large interest in investigating how multilingual LLMs organize and share knowledge between languages internally. A key question is whether LLMs “think” in English, or whether they rely on English-centric representations or reasoning steps even when operating in other languages. There is evidence that multilingual models often map other languages into an implicit English latent space. For example, [36] observes that when posed a math question in another language, a language-aligned LLM will typically generate an English chain-of-thought before producing the final answer. [15] uses bilingual sentence pairs to pull representations of translations closer together in vector space. Other works analyze the intermediate embeddings of transformer across layers when input with prompt in English or low-resource language, and found distinct phases of operation, where middle layers perform reasoning in a latent space steer toward the dominant language (English) [30, 26]. This suggests the model’s internal reasoning happens in English by default.
Figure 2: Loss curves over the training process for Native CoT Training (Left) and our approach, English-Pivoted CoT Training (right), normalized with exponential moving average with a smoothing weight of 0.95. Our method shows lower initial reasoning loss and slower decline, indicating effective separation of English reasoning from target-language responses.
Model editing techniques have been applied to leverage these insights to boost performance of multilingual LLMs for specific tasks. For example, [24] leverages model merging techniques to combine between a LLM with expertize in a low-resource language, with another LLM with expertize in math for English. [2] does not combine weights, but only swaps layers between the 2 models, and still achieves strong performance.
These alignment efforts all seek to improve cross-lingual generalization, allowing knowledge learned in one language to transfer to others. Our work is informed by these studies but takes a distinct direction. In contrast to prior work that often relies on external models or separate bilingual modules, our approach aligns the reasoning process within the model’s own representations. This allows us to transfer reasoning skills to low-resource languages without retraining from scratch, advancing the understanding of how multilingual LLMs can maintain consistent reasoning across diverse languages.
3 Bridging Test-time Scaling for Low-resource Languages
Given the recent emergence and popularity of post-training methods for language models, especially their demonstrated effectiveness in improving reasoning task accuracy, this paper explores extending test-time scaling to low-resource languages. Such languages are typically underrepresented in both pre-training and post-training corpora, limiting model performance. Our approach, named English-Pivoted CoT Training, leverages the model’s robust reasoning capabilities developed predominantly in English, by explicitly constraining the intermediate chain-of-thought reasoning steps to English, while maintaining inputs and final responses in the target low-resource language.
Figure 3: Parameter updates (mean absolute differences) of Native CoT Training (Left) and our approach, English-Pivoted CoT Training (right) for r1-distill-Llama-8B. English-Pivoted CoT Training makes fewer changes to the base model, and focuses more on language understanding and generation layers (red boxes).
Formally, let $x$ denote an input problem expressed in the target (low-resource) language, and let $y$ represent the final answer also in the target language. We denote the chain-of-thought reasoning trace as $C$, which is constrained to be in English. The rationale behind this design is that models typically have been extensively trained in English for reasoning tasks, and we hypothesize that their latent spaces are more robustly aligned with English reasoning processes. Our method is designed to learn the conditional probability distribution:
$P_{\theta}(C,y|x)=P_{\theta}(C|x)P_{\theta}(y|x,C)$ (1)
where $\theta$ represents the parameters of the language model, $C=(s_{1},\dots,s_{n})$ is the chain-of-thought, comprising natural language reasoning steps $s_{i}$. Following recent works [32, 20], we explicitly separate $C$ and $y$ by a special token (e.g. “</think>”) to mark the transition between the reasoning trace and the final answer.
Given a training dataset $D=\{(x,C,y)\}$ consisting of tuples where:
  • $x\in S_{T}$: input in the target low-resource language.
  • $C\in S_{en}$: reasoning trace in English.
  • $y\in S_{T}$: final answer in the target low-resource language.
we optimize the following objective during training:
$\mathcal{L}(\theta)$ $=\sum_{x,C,y\in D}\log P_{\theta}(C,y|x)$ (2)
$=\sum_{x,C,y\in D}[\log P_{\theta}(C|x)+\log P_{\theta}(y|x,\hat{C})]$ (3)
$=\sum_{x,C\in D}\log P_{\theta}(C|x)+\sum_{x,y\in D}\log P_{\theta}(y|x,\hat{C})$ (4)
$=\alpha\sum_{x,C\in D}\log P_{\theta}(C|x)+\beta\sum_{x,y\in D}\log P_{\theta}(y|x,\hat{C})$ (5)
where $\log P_{\theta}(C|x)$ optimizes the model to generate English reasoning traces given input in different languages, and $\log P_{\theta}(y|x,\hat{C})$ with $\hat{C}=\theta(x)$ pushes the model to generate final response based on both the input (in target language) and the generated reasoning trace. Additionally, each objective can be balanced with hyperparameter weights $\alpha$ and $\beta$. In this initial work, we set both the hyperparameters $\alpha$ and $\beta$ to 1 for simplicity.
Effectively, this training approach allows the model to “reason” in English (through forcing the ground-truth chain-of-thought to be in English and the loss function $\log P_{\theta}(C|x)$), a language where it has been extensively trained for reasoning. In other words, this enhances the representation alignment between input prompts in different languages, where they will lead to the same traces of reasoning. The training method also allows the model to understand and provide responses in the low-resource target language. This not only boosts performance on low-resource languages but also simplifies the training process by reducing the need for large amount of data in those languages. In Figure 2, we visualize the changes of the two loss components in Equation (5) over the first training epoch. The figure indicates that our English-Pivoted CoT training makes it easier for the model to maintain reasoning in English, as evidenced by the lower initial reasoning loss and slower subsequent decrease. The increasing gap between reasoning and response losses shows our method effectively separates English reasoning from target-language generation, enabling better understanding of the low-resource language. In contrast, Native CoT starts with higher initial reasoning loss that decreases rapidly, resulting in a smaller gap between the two losses, indicating less separation between language understanding and reasoning processes.
This effect can also be seen in Figure 3, comparing the parameter update patterns of Native CoT training (multilingual reasoning fine-tune with both CoT traces and answers in target language) and our proposed English-Pivoted CoT training for r1-distill-Llama-8B. The figure illustrates key differences in adaptation dynamics. Firstly, considering the absolute magnitude of parameter changes, Native CoT introduces larger updates throughout the model (approximately 1.3 times larger), suggesting a more significant departure from the baseline model’s parameters, which can lead to substantially different reasoning behavior. Secondly, examining the relative changes across model layers, English-Pivoted CoT also concentrates updates in the first and last few layers, particularly within attention-related matrices (highlighted in red boxes). These layers have been shown in prior research to be in charge of language understanding and generation tasks [26, 30]. Consequently, our method strategically targets these layers, enabling the model to effectively understand and generate final responses in the low-resource language without deviating from the original model’s reasoning capability.
4 Experiments
4.1 Experiment settings
Baselines. We select r1-distill-Llama-8B [4] and DeepHermes-3-Llama-3-8B [25] due to their state-of-the-art performance and open-source availability. We perform fine-tuning on these models following our proposed approach (denote English-Pivoted CoT Training (OURS)). We compare the performance of our method against several existing multilingual adaptation techniques, including multilingual reasoning fine-tuning (denote Native CoT Training), model merging and layer swap.
Given the limited availability of reasoning data in low-resource languages, a practical approach is leveraging machine translation. Specifically, we use the NLLB translation model to translate existing reasoning datasets into our target languages Bespoke [12] into Irish, pensez [7] into French. For Chinese, a high-resource language, a corpus is available, named congliu [16]. We sample the amount of data in target language to be 10,000 samples, and also add in 10.000 English samples from the same source to prevent forgetting. In our approach, reasoning traces originally in target languages are translated into English for training purposes. While the amount of data we leverage for adaptation to low-resource languages is relatively small (e.g., compared to 800,000 samples used in [4]), the evaluation results presented in Section 4.2 illustrate a clear improvement for our method, highlighting its effectiveness. We note that all evaluation datasets are created or manually verified by humans, ensuring accuracy and quality.
Training setup. Our training implementation employs HuggingFace Transformers and DeepSpeed. To manage memory constraints efficiently, we set the maximum input sequence length to 16,384 tokens, reduced from the original 32,768. Models are trained using the AdamW optimizer [17] for 3 epochs, with a maximum learning rate of $1\times 10^{-5}$. Training is distributed across two Nvidia A100 GPUs (80GB each), with a total batch size of 24.
Benchmarks. We evaluate the reasoning capabilities of all models on standard English reasoning benchmarks, including merican Invitational Mathematics Examination 2024 - AIME2024 [28] (challenging Math Olympiad-level problems) and MGSM [23] (high school-level math problems). Additionally, we use MGSM for evaluating performance in French and Chinese (as it is a multilingual benchmark dataset), supplemented by the MATH-hard dataset [18] (a translation of the MATH500 dataset [8], keeping only level-5 difficulty competition math problems), for French, consistent with the French LLM leaderboard [18].
On the other hand, for the Irish language, due to the lack of existing evaluation datasets, we introduce two new evaluation benchmarks:
Table 1: Performance comparison of our English-Pivoted CoT Training approach against baseline methods across English (en) and Irish (ga) reasoning benchmarks. Bold scores indicate the best performance per benchmark.
Model Setting AIME2024 (en) AIME2024 (ga) LC2024 (ga)
r1-distill-Llama-8B Base 43.33 6.67 63.64
r1-distill-Llama-8B Native CoT Training [1, 13] 48.33 21.67 37.14
r1-distill-Llama-8B TIES-merging [31, 24] 0.00 1.67 20.00
r1-distill-Llama-8B Layer Swap [2] 0.00 1.67 0.00
r1-distill-Llama-8B English-Pivoted CoT Training (OURS) 53.33 35.00 73.33
DeepHermes-3-Llama-3-8B Base 1.67 6.67 52.73
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 1.67 8.33 40.01
DeepHermes-3-Llama-3-8B TIES-merging [31, 24] 0.00 0.00 1.82
DeepHermes-3-Llama-3-8B Layer Swap [2] 0.00 0.00 12.73
DeepHermes-3-Llama-3-8B English-Pivoted CoT Training (OURS) 5.00 10.00 54.55
Table 2: Comparisons of low-resource language understanding on LC2024, where the exam is split into 2 parts: concepts & skills and contexts & applications. The latter part has additional contextual and real-world descriptions, requiring a higher level of Irish language understanding to comprehend the input. Bold scores indicate the best performance per benchmark.
Model Setting LC2024 - concepts & skills LC2024 - contexts & applications
r1-distill-Llama-8B Base 84.85 31.82
r1-distill-Llama-8B Native CoT Training [1, 13] 47.50 31.82
r1-distill-Llama-8B TIES-merging [31, 24] 27.27 9.09
r1-distill-Llama-8B Layer Swap [2] 0.00 0.00
r1-distill-Llama-8B English-Pivoted CoT Training (OURS) 82.82 59.09
DeepHermes-3-Llama-3-8B Base 57.58 45.45
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 57.58 13.64
DeepHermes-3-Llama-3-8B TIES-merging [31, 24] 3.03 0.00
DeepHermes-3-Llama-3-8B Layer Swap [2] 9.10 18.18
DeepHermes-3-Llama-3-8B English-Pivoted CoT Training (OURS) 57.58 50.00
  • Irish version of AIME2024: translated and verified by two Irish speakers.
  • Leaving Certificate 2024 Math exam - Higher Level (LC2024): Derived from official Irish Leaving Certificate examination materials. [^1] https://www.examinations.ie/archive/exampapers/2024/LC003ALP100IV.pdf [^2] https://www.examinations.ie/archive/exampapers/2024/LC003ALP200IV.pdf We extract individual questions that have concrete answers (e.g., excluding proof-based questions) and do not require additional modalities beyond text (e.g., geometric diagrams). The questions are formatted in LaTeX, resulting in a total of 55 unique samples.
During evaluation, we follow the training configuration by limiting generation outputs to a maximum length of 16,384 tokens. We adopt evaluation hyperparameters consistent with each benchmark’s original protocols unless otherwise specified. When not explicitly stated, we use a decoding temperature of 0.6 and top-k sampling with $k=0.95$.
Figure 4: Representation retrieval accuraccy between Left: questions, and Right: questions and reasoning traces of the same questions in different languages.
Table 3: Ablation study of fine-tuning on French - a medium resource language, and benchmark across English and Fench reasoning datasets. Bold scores indicate the best performance per benchmark.
Model Setting AIME2024 (en) MSGM (en) MATH-hard (fr) MSGM (fr)
r1-distill-Llama-8B Base 43.33 79.6 49.74 54.8
r1-distill-Llama-8B Native CoT Training [1, 13] 45.00 82.0 31.90 61.2
r1-distill-Llama-8B English-Pivoted CoT Training (OURS) 50.00 89.6 70.01 83.2
DeepHermes-3-Llama-3-8B Base 1.67 38.8 3.69 21.6
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 3.33 70.8 6.26 44.8
DeepHermes-3-Llama-3-8B English-Pivoted CoT Training (OURS) 8.33 76.8 32.27 77.2
Table 4: Ablation study of fine-tuning on Chinese - a high resource language, where baselines have high level of understanding and tend to output CoT traces in Chinese. Bold scores indicate the best performance per benchmark.
Model Setting AIME2024 (en) MSGM (en) MSGM (zh)
r1-distill-Llama-8B Base 43.33 79.6 65.2
r1-distill-Llama-8B Native CoT Training [1, 13] 46.33 92.4 81.6
r1-distill-Llama-8B English-Pivoted CoT Training (OURS) 56.67 92.4 70.0
DeepHermes-3-Llama-3-8B Base 1.67 85.2 50.8
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 11.67 83.2 61.6
DeepHermes-3-Llama-3-8B English-Pivoted CoT Training (OURS) 13.33 89.6 62.4
4.2 Results and analysis
Reasoning capability can be effectively transferred to solve problems in low-resource language. Table 1 presents the performance of models trained with our approach (English-Pivoted CoT Training), compared to other baselines on English and Irish benchmarks. Our method demonstrates a strong improvement over the baselines. On the Irish version of AIME2024, while all baselines perform poorly, our approach obtains a clear gap of upto 28.33%, in the case of the r1-distill-Llama-8B baseline. More specifically, we achieve accuracies of 35.00% and 10.00% when fine-tuned on r1-distill-Llama-8B and DeepHermes-3-Llama-3-8B, respectively, matching performance on the English version. On LC2024, which comprises relatively less challenging math problems than AIME2024, although the r1-distill-Llama-8B baseline achieves an acceptable accuracy of 63.64%, our approach surpasses it by a 10% margin (73.33% vs. 63.64%). Further inspection shows that both baseline models generate all chain-of-thoughts and final answers in English, regardless of prompting in which languages. This default behavior aligns with the prompting approaches in [21, 23], where they prompt engineering the model to reason in English to obtain higher performances.
Furthermore, the TIES-merging and Layer Swap approaches fail and even unable to correctly answer any of the questions on AIME2024, likely due to significant parameter discrepancies when merging the base model trained on English reasoning tasks with a model trained on general instruction tasks in the low-resource language. These differences likely cause interference between subsets of merged parameters. Another interesting observation is that Native CoT Training, training entirely in the target language for both the chain-of-thought and final answer improved performance on AIME2024 but degraded it on LC2024 (63.64% to 37.14% for r1-distill-Llama-8B), potentially due to reduced reasoning capability when forced to reason in a low-resource language.
Improved language understanding capability. To further investigate whether our approach also enhances language understanding or simply overfits to mathematical phrases in the target language, we split LC2024 into its two original sections: Concepts and Skills and Contexts and Applications. The latter, comprised of more contextual and real-world application descriptions, requires deeper language understanding in Irish. As shown in Table 2, Native CoT training significantly reduces the model’s reasoning capability, as evidenced by the drop from 45.45 to 13.64 on Concepts and Skills for DeepHermes-3-Llama-3-8B, while maintaining the same level of performance for r1-distill-Llama-8B. This supports the idea that forcing the model to reason in the low-resource language affects its ability to generate effective reasoning traces. In contrast, our approach yields gains substantially in Contexts and Applications split — indicating that not only does English-Pivoted CoT Training enable effective reasoning in low-resource languages, but it also improves overall language understanding.
Figure 5: Changes in correction rate from baseline model (r1-distill-Llama-8B) when fine-tuned with our proposed approach across 3 benchmark datasets on 3 languages, from low (Irish) to medium (French) to high (Chinese) resource.
Diverged reasoning traces when forced to reason in different languages. Furthermore, we analyze the model’s representations across languages by computing and comparing the average token representations for both the input alone and the input plus the generated reasoning trace on AIME2024. We measure the retrieval accuracy for matching representations across languages with either questions or questions concatenating with reasoning traces on AIME2024. The retrieval task involves finding the closest English sample given an Irish sample representation, using average token embeddings at any layer. Figure 4 visualizes the accuracies across layers. While models trained solely in the target language (Native CoT Training) show strong alignment for questions (even upto 100% at middle layers), their alignment significantly drops (to highest of around 35% at 2 layers) when a chain-of-thought is included — highlighting discrepancies when the model is forced to reason in a different language than the final output.
This also suggests that question alignment [36, 14, 9], either through parallel sample fine-tuning or prompting, is not enough. Figure 4 shows that while the alignment between questions in different languages is inherently high, this can still lead to really diverged reasonings and answers. Figure 1 depicts an example, where the whole embedding of the questions are similar, a small difference (e.g., mixing between independence and dependence) can lead to a totally different interpretation.
In contrast, our approach - English-Pivoted CoT Training, which trains the model to understand the target language while reasoning in English, achieves the strongest alignment between the reasoning traces (almost close to 100%), demonstrating a more consistent internal representation.
Generalizability to medium- and high-resource languages. We apply our proposed approach to Chinese and French languages, which can be considered medium to high-resource languages [33, 3]. The results, present in Table 3 and Table 4, indicate that our approach are generalizable to other languages. First, the baseline models perform better compared to on the extremely low-resource language, Irish, with a score of 54.8% and 65.2% accuracy on MSGM (French and Chinese versions, respectively) for r1-distill-Llama-8B. Nevertheless, we see a sustainable improvement on French when leveraging our approach, of 83.2% to 54.8% on MGSM (French version). However, on Chinese, as the model already performs well, our approach does not lead to much improvement. Furthermore, qualitative analysis of the reasoning traces of the base models shows that the models actually perform reasoning with CoT in Chinese, compared to English, when prompted with Chinese problems versus problems in low and medium-resource languages. Therefore, forcing the model to reason in English can create an interference. This also explains why further training on target-language reasoning traces (Native CoT Training) can improve the performance (81.6% compared to 70.0% of OURS and 65.2% of baseline).
Figure 5 shows the correction rate of fine-tuned model with our approach, compared to the baseline r1-distill-Llama-8B. More specifically, the percentage of the new model corrects an incorrect answer by the base model and vice versa. We can see that on MSGM (zh), the correct $\rightarrow$ incorrect rate is the highest, highlighting conflicts and diverged solutions by forcing the model to think in English for Chinese problems when they already have a tendency to think in Chinese for Chinese problems. Nevertheless, our method still achieves an average improvement of 14.30% across the 3 resource regimes (low, medium, and high-resources), demonstrating its generalizability.
5 Conclusion
In this work, we propose a novel training approach, named English-Pivoted CoT Training, to effectively transfer the reasoning capabilities of large language models to low-resource languages. Through training to explicitly aligning reasoning processes across languages by forcing CoT traces to be in the dominant language, our approach achieves significant performance gains (up to 28.33%) over existing multilingual reasoning techniques, and is generalizable to other resource regimes (medium and high-resource languages). Furthermore, our analysis provide insights into multilingual LLM behavior, particularly the benefit of separating language understanding from reasoning. By contributing the first-ever Irish mathematical reasoning benchmark (LC2024), we also aim to support future research in multilingual reasoning. Future directions include exploring the generalizability of our approach to other low-resource languages and tasks beyond mathematics, as well as investigating strategies for further improving cross-lingual reasoning capabilities.
{ack}
This research work has emanated from research conducted with financial support from Science Foundation Ireland under Grant 12/RC/2289-P2 and 18/CRT/6223.